-(5y+3)-(-4y-6)=-1
Answer by ankor@dixie-net.com(22740)

-(5y+3)-(-4y-6)=-1
Remove brackets (neg outside the brackets, changes the signs inside the brackets)
-5y - 3 + 4y + 6 = -1
-5y + 4y + 3 = -1
-y = -1 - 3
-y = -4
y = +4; multiply equation by -1 to get a positive y
;
;
Check solution in original equation
-(5(4) + 3) - (-4(4) - 6) = -1
;
-(23) - (-22) = -1
:
-23 + 22 = -1; confirms our solution